JustALilCat:

Which equation results from adding the equations in this system? x + 6 y = 9. Negative x + 2 y = negative 15. A) 8 y = negative 6 B) 8 y = 6 C) 4 y = 24 D) 4 y = negative 6

JustALilCat:

I'm horrible in math ;-;

iYuko:

So sorry if im bad at explaining this but its not difficult! You're basically just going to add each variable together. x + 6y = 9 -x+2y= -15 So the X is going to cancel out so were left with: 6y=9 and 2y=-15 Now we just add 8y=-6 Does that make sense ?

JustALilCat:

@iyuko wrote:
So sorry if im bad at explaining this but its not difficult! You're basically just going to add each variable together. x + 6y = 9 -x+2y= -15 So the X is going to cancel out so were left with: 6y=9 and 2y=-15 Now we just add 8y=-6 Does that make sense ?
No, I did not understand any of that, sadly

JustALilCat:

But thank you for the help anyway, @iyuko

JustALilCat:

@justalilcat wrote:
But thank you for the help anyway, @iyuko
Also, you have a very cool username

iYuko:

@justalilcat wrote:
@justalilcat wrote:
But thank you for the help anyway, @iyuko
Also, you have a very cool username
Thx!

iYuko:

@justalilcat wrote:
@iyuko wrote:
So sorry if im bad at explaining this but its not difficult! You're basically just going to add each variable together. x + 6y = 9 -x+2y= -15 So the X is going to cancel out so were left with: 6y=9 and 2y=-15 Now we just add 8y=-6 Does that make sense ?
No, I did not understand any of that, sadly
So like variables can be added together x + (-x) = 0 so it's canceled out so were left with: 6y=9 and 2y=-15 6y+2y= 8y and 9 + (-15) = -6 So your final answer is A) 8 y = negative 6

JustALilCat:

@iyuko wrote:
@justalilcat wrote:
@iyuko wrote:
So sorry if im bad at explaining this but its not difficult! You're basically just going to add each variable together. x + 6y = 9 -x+2y= -15 So the X is going to cancel out so were left with: 6y=9 and 2y=-15 Now we just add 8y=-6 Does that make sense ?
No, I did not understand any of that, sadly
So like variables can be added together x + (-x) = 0 so it's canceled out so were left with: 6y=9 and 2y=-15 6y+2y= 8y and 9 + (-15) = -6 So your final answer is A) 8 y = negative 6
Ohhhh, now I think I get it, I basically take the letter from the number, and then solve from there?

iYuko:

Like terms can be added together x and -x are like terms just like 6y and 2y. You can add them together like normal and just carry the y...so it makes 8y. It's really simple so don't overthink it!